OpenStudy (anonymous):

MEDAL AND FAN AWARDED Solve for x: x - 2 > 2x + 12. x > 14 x < 14 x > -14 x < -14 Evaluate 10 + 8 ÷ 2 – 4. 10 5 6 14

OpenStudy (anonymous):

@133

OpenStudy (anonymous):

@AMYCARTER

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

okay let's solve the first one. x−2>2x+12 Step 1: Subtract 2x from both sides. x−2−2x>2x+12−2x −x−2>12 Step 2: Add 2 to both sides. −x−2+2>12+2 −x>14 Step 3: Divide both sides by -1. −x _____ −1 > 14 −1 x<−14 Answer: x<−14

OpenStudy (anonymous):

Alright, I think I understand. Thanks!

OpenStudy (anonymous):

the second one is 5. 10+8 is 18, 18 divided by 2 is 9 and 9 - 4 is 5

OpenStudy (133):

The first one is x<-14 and the second one is 10

OpenStudy (anonymous):

Wouldn't you start off with dividing? As PEMDAS puts the functions in order?

OpenStudy (133):

@AMYCARTER Did you do PEMDAS??

OpenStudy (anonymous):

It would be division first, then addition, and subtraction making it 10

OpenStudy (133):

Because you would do 10 + 8 ÷ 2 – 4 8 divided by 2=4 10+4-4 10+4=14 14-4=10 So it would be 10 @CoinBank77 why would you ask the question if you knew the answer??

OpenStudy (anonymous):

I figured it out while I was waiting for replies, do you think you could help me with a few more?

OpenStudy (133):

Sure but I've only got 6 min to help

OpenStudy (anonymous):

Alex stores fertilizer for his farm in a rectangular box. The length of the box is 6 feet, width is 3 inches, and height is 9 inches. Each bag of fertilizer contains 0.225 cubic foot of fertilizer. How many bags of fertilizer will fill the box completely? [1 foot = 12 inches] Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

@133 alright, it shouldn't be much longer.

OpenStudy (133):

So you would have to figure out the area of the box.

OpenStudy (anonymous):

72x3x9?

OpenStudy (133):

2(wl+hl+hw) you plug it all in it would equal 198

OpenStudy (anonymous):

The answer is 198 or that's the total area?

OpenStudy (133):

That's the total area now we need to do 198*0.225=??

OpenStudy (anonymous):

44.55

OpenStudy (anonymous):

That's the answer? No rounding or anything else?

OpenStudy (133):

Yes and no but I have to go now

OpenStudy (anonymous):

Alright, thanks for the help!

OpenStudy (133):

No problem